Find the corresponding area between 𝑧 = 0 and each of the following.

1. z = 0.85

2. z = 1.27

3. z = 2.86

4. z = βˆ’1.05

5. z = βˆ’2.96


1
Expert's answer
2021-05-02T08:19:28-0400

1.


"P(0<Z<0.85)=0.8023-0.5=0.3023"

2.


"P(0<Z<1.27)=0.8980-0.5=0.3980"

3.


"P(0<Z<2.86)=0.9979-0.5=0.4979"

4.


"P(-1.05<Z<0)=0.5-0.1469=0.3531"

5.


"P(-2.96<Z<0)=0.5-0.0015=0.4985"